Please solve,wordproblemsare so difficult. :) Thank you!
The trip upstream from A to B takes 4.5 hrs, and the trip back downstream takes 2 hrs. Its 18 miles from A toB. What is the speed of the current and what is the speed of the boat in still water?
Thank you!

Answer by scott8148(6628)

let b=boat, c=current ... speed upstream=18/4.5=4 ... speed downstream=18/2=9

b+c=9 ... b-c=4 ... b+c+(b-c)=9+4 ... 2b=13 ... b=6.5 ... c=9-b ... c=2.5